The pth term of an AP is (3p-1)/6. The sum of the first n terms of the AP is?
Answer by stanbon(75887)

The pth term of an AP is (3p-1)/6. The sum of the first n terms of the AP is?
------
Sum Formula: S(n) = (n/2)(a(1)+a(n)
-------------------
a(1) = (3-1)/6 = 1/2
a(n) = (3n-1)/6
-----
S(n) = (n/2)[(1/2)+(3n-1)/6]
----
= (n/2)[(3+3n-1)/6]
====
= (n/2)(3n+2)/6
----
= (3n^2+2n)/12
=====================
Cheers,
Stan H.
---------------------